OpenStudy (anonymous):

x - 5 less than or equal to 8

OpenStudy (anonymous):

Please help...

jimthompson5910 (jim_thompson5910):

add 5 to both sides \[\Large x -5 \le 8\] \[\Large x -5+5 \le 8+5\] \[\Large x +0 \le 13\] \[\Large x \le 13\]